Paul orders a pizza. Chef carl randomly chooses 2 different toppings to put on the pizza from the following: pepperoni, onion, sausage, mushrooms, and anchovies. if paul will not eat pizza with mushrooms, determine the probability that paul will not eat the pizza chef carl has made.
Found 2 solutions by usyim88hk, JWG:
Answer by usyim88hk(157)

We can do it with a shortcut call "combination", but I will use the long way, call "list them all out!"
Different pizzas that Paul can order:
1) Peperoni, onion
2) Peperoni, sausage
3) Peperoni, mushrooms
4) Peperoni, anchovies
5) onion, sausage
6) onion, mushrooms
7) onion, anchovies
8) sausage, mushrooms
9) sausage, anchovies
10) mushrooms, anchovies
So now we can see that 4 out of 10 pizzas have mushrooms on them, so the probability will be 40%

Answer by JWG(21)

This is an A or B Probability Question. We want to know the probability of picking mushroom as a first topping or a second topping.
Probability of picking mushroom as the first topping is 2/5 (We'll call this A)
Probability of picking mushroom as the second topping is 1/4 (We'll call this B)
The Probability of A or B Formula is A + B - A x B
A + B = 2/5 + 1/4 = 4/20 + 5/20 = 9/20
A x B = 1/5 x 1/4 = 1/20
A + B - A x B = 9/20 - 1/20 = 8/20 = 4/10 = 2/5